OpenStudy (anonymous):

write the equations in standard form of two quadratic functions that have the same vertex but open in different directions.

OpenStudy (cwrw238):

first write in vertex form (x-1)^2 + 2 -(x-1)^2 +2

OpenStudy (cwrw238):

the -1 in the brackets gives the x coordinate of the vertex The negative in front of the brackets in second equation flips the whole graph in the y-direction

OpenStudy (cwrw238):

so we now expand the eqautions to standard form; x^2 - 2x + 3 = 0 and -x^2 +2x +1 = 0

OpenStudy (anonymous):

Thank you so much for your help! I completely understand how to do this, except one part: how did you get the constant in the second function in standard form as 1?

OpenStudy (cwrw238):

-(x-1)^2 + 2 = -(x^2-2x + 1) +2 = -x^2 +2x -1 +2 = -x^2 +2x + 1

OpenStudy (cwrw238):

'm having difficulty with this website - i cant post a large message more that a few lines so i have to squash it up!

OpenStudy (anonymous):

It's no problem really, thank you so much!!